Actuallly, if my memory of statistics is correct, they would have roughly a 24% chance of winning 2 games. You have to multiply the chance of winning by itself, then multiply that by the number of possibilities for winning 2 games. Odds of winning 2 games = .2 * .2 = .04 (4%) Winning 2 Game Possibilities: 1,2 - 1,3 - 1,4 - 2,3 - 2,4 - 3,4 = 6 possibilites .04 * 6 = .24 = 24% chance of winning 2 games (if they have a 20% chance of winning each game individually). Anyhow, I agree, a series sweep is probably wishful thinking there. I think 2 or 3 games is more realistic. I also believe that we can take 2 from the Dodgers, not just 1.